Determine the molecular formula of an oxide of sulphur containing 50% sulphur

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear pallavi

percentage of sulpher =50%

percentage of oxygen = 100-50 =50%

 number of molecule of sulpher = 50/32=1.5625

 number of  molecules of oxygen = 50/16 =3.125

 

 so simple ration of molecules are = S/O = 1/2

so emperical formula = SO2

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you the answer and detailed
solution very  quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin